OpenStudy (anonymous):

i am lost and i dont know where to start, i know this is calculus 1 but this is calculus 2. the section is Applications of Definite Integrals and i am trying to find the lengths of the curves. The question is x=(y^3/3)+1/(4y) from y=1 to y=3

OpenStudy (anonymous):

length or area under the curve?

OpenStudy (anonymous):

Integrate the equation with respect to y and use y=3 and y=1 for upper and lower limits respectively. You'll have area under the graph

OpenStudy (anonymous):

i understand that but i need help setting it up for the integral, after that i can solve it,

OpenStudy (anonymous):

Arc length?

OpenStudy (anonymous):

no it just says lengths of the curves, between 1-3

OpenStudy (anonymous):

Yeah, that's usually solving for arc lengths, you'll have to find dx/dy first, here the "formula" if you will... \[L = \int\limits_{a}^{b} \sqrt{1+(\frac{ dx }{ dy }})^2dy\] as in your case, see what you get.

OpenStudy (anonymous):

well than is my problem, i dont know how to get it set up for the intrgral, to get (y^3/3)+1/(4y) to be ready to integrate,

OpenStudy (anonymous):

Do one step at a time, as it shows, so find dx/dy first and then square it, etc.

OpenStudy (anonymous):

so square (y^3/3+1/(4y))^2 like that?

OpenStudy (anonymous):

But you have to find dx/dy first

OpenStudy (anonymous):

ok then thats where i am lost now

OpenStudy (anonymous):

so then would it be y^2-1/2y^2

OpenStudy (anonymous):

\[\frac{ dx }{ dy } = y^2-\frac{ 1 }{ 4y^2 }\]

OpenStudy (anonymous):

Now you have \[1+(y^2-\frac{ 1 }{ 4y^2 })^2\]

OpenStudy (anonymous):

Work with that and squareroot it and then you can integrate it easily :P, I'll bbs, so keep working on it!

OpenStudy (anonymous):

ok so far i am getting that

OpenStudy (anonymous):

Actually, \[\sqrt{1+(\frac{ dx }{ dy }})^2 = y^2+\frac{ 1 }{ 4y^2 }\] so you just have to integrate this from 1 to 3

OpenStudy (anonymous):

That's a very easy integral!

OpenStudy (anonymous):

\[\int\limits_{1}^{3} y^2+\frac{ 1 }{ 4y^2 } dy\]

OpenStudy (anonymous):

yup that is what i have, thank you, integral part i can do, the setting up for it throws me off

OpenStudy (anonymous):

No problem, remember take it step by step :)